# Exploring Performance–Complexity Trade-Offs in Sound Event Detection Models

### v1

@@ Abstract
Previous research on low-complexity audio classification has leveraged efficient convolutional neural network (CNN) architectures adapted from the vision domain [6, 7, 8] or introduced custom-designed efficient CNNs [9, 10, 11, 12, 13]. However, most of these models are optimized for clip-wise predictions, i.e., predicting a class label for an entire given audio clip. In contrast, the goal in SED is to identify specific acoustic events in a recording, including begin and end times, which implies a need for frame-wise predictions [14].
Unlike clip-wise tasks, where local features from the CNN are typically aggregated using global pooling, SED commonly employs a sequence model on top of the CNN to capture long-term temporal dependencies. The most common choice for the sequence model is a recurrent neural network (RNN), resulting in widely used CRNNs [15, 16, 17, 18]. However, the optimal sequence model on top of a CNN for achieving the best performance–complexity trade-off remains an open question—one that we address in this paper. The closest to our work is LCSED [19], which proposes a lightweight SED model composed of a hybrid CNN and a dual-attention module as a more efficient variant to the traditionally used CRNN [15, 16, 17, 18].
This work is more general, as we propose a series of models that combine a CNN backbone with various sequence models, and explore their performance-complexity trade-offs. We build on the MobileNetV3 [20] architecture as the CNN backbone, as it is a well-established low-complexity architecture that has shown excellent performance for large-scale audio tagging [8], a clip-wise task. To get the most meaningful results, we evaluate on the largest publicly-available SED dataset - AudioSet Strong [21] - and use the state-of-the-art training pipeline [4], which includes knowledge distillation [22] from large transformer models. Our experiments aim to answer the following key questions:
@@ II Architectures
where and represent a linear projection aligning embeddings with the dimension of the sequence model . If no sequence model is used, and the linear projection (, ) are identity functions. The output of is then transformed by a position-wise linear layer, parameterized by and , to generate frame-level predictions , with dimension corresponding to the number of classes in the task (e.g., 447 for the training set of AudioSet Strong).
@@ II-A Frame-wise Convolutional Backbone
As the convolutional backbone, we use MobileNetV3 [20]. To enable frame-wise instead of clip-wise predictions, we remove the global pooling and the global classification heads. The outputs of this network are then three-dimensional embeddings of shape . Compared to the original MobileNet, we increase some convolutional strides along the frequency axis from 1 to 2 while decreasing others along the time axis from 2 to 1. This changes the shape of from to . Therefore, each of the 250 time frames in corresponds to a 40-millisecond segment of the original 10-second audio clip. Since only stride values are modified, the adapted model remains compatible with the pre-trained ImageNet [23] checkpoint of the original counterpart. We refer to this adapted model as frame-wise MobileNet (fmn).
For the experiments that require it, we scale the frame-wise MobileNets following the method described in [8], maintaining the number of layers while adjusting the model’s width (i.e., number of channels) using a scaling factor . We consider models with which we name fmn04, fmn06, fmn10, fmn20 and fmn30, respectively.
@@ III-B Training Configuration
We initialize the backbone CNNs with ImageNet [23] pre-trained weights and train them on AudioSet Weak [34], following the knowledge distillation routine from [8]. However, we find that extending the training to 300 epochs improves performance, achieving mean average precisions of .420, .437, .458, .474, and .475 for fmn04, fmn06, fmn10, fmn20 and fmn30, respectively.
The full models (backbone + sequence model) are trained on AudioSet Strong for 120 epochs using an Adam optimizer with no weight decay, a cosine learning rate schedule with 5,000 warmup steps, and a batch size of 256.
@@ IV-A Performance-Complexity Trade-Off of Sequence Models
Figure 1 shows that all sequence models improve the performance of fmn10, even with the smallest hidden dimension. However, only models incorporating TF, BiGRU, ATT, and HYBRID achieve performance comparable to or surpassing fmn20 while maintaining lower parameter counts, lower MACs, and increased throughput, resulting in a more favorable performance-complexity trade-off. This shows that adding suitable sequence models on top of convolutional backbones scales more efficiently than simply scaling up the convolution backbone itself. In contrast, MAMBA and TCN fail to reach the performance level of fmn20 at any hidden dimension size.
The choice of the model offering the best performance-complexity trade-off depends on the complexity metric of interest. The top choices for parameters, throughput, and MACs are HYBRID, TF, and BiGRU, respectively. For example, fmn10 using TF and a hidden dimension of 256 (fmn10+TF:256) matches the performance of fmn20 while achieving 1.76 times its throughput and requiring only 37.6 of its total parameters. Similarly, fmn10 with HYBRID at a hidden dimension of 128 (fmn10+HYBRID:128) has fewer parameters than fmn10 without a sequence model444 If the down-projection, sequence model, and linear head together require fewer parameters or MACs than a single large linear layer mapping embeddings to target classes, adding a small sequence model can be more efficient than using no sequence model with respect to these metrics. yet improves the PSDS1 score from 42.28 to 43.45. However, fmn10+HYBRID:128 exhibits lower throughput compared to fmn10+TF:256. In terms of MACs, BiGRU is the most effective555BiGRUs involve multiple operations, such as gating, point-wise products, and point-wise summations, which do not count as MACs., but due to its sequential nature, it lags behind TF and HYBRID in terms of throughput. We conclude that TF is the most balanced choice with strong results on all three complexity measures. In particular, fmn10+TF:256 strikes an excellent trade-off between performance gain and minimal added complexity, making it the preferred model for further experiments.
@@ IV-B Scaling the Backbone and Sequence Model
Three key conclusions can be drawn from Figure 2: 1) TF and HYBRID consistently outperform fmns without sequence models across all scaling factors; 2) HYBRID performs best in the lower complexity range, with fmn04+HYBRID and fmn06+HYBRID surpassing their TF and BiGRU counterparts; and 3) TF and HYBRID exhibit greater stability than BiGRU, as the performance drop of BiGRU from to is a consistent trend across multiple independent runs, not an outlier.
@@ IV-C Impact of KD and CNN Backbone Fine-Tuning
Model | KD | Frozen Backbone | PSDS1 |
@@
Table I presents the impact of Knowledge Distillation and fine-tuning the pre-trained convolutional backbone on AudioSet Strong. Regardless of whether a sequence model (TF:256 on top of fmn10) is used, KD significantly enhances performance. Similarly, fine-tuning the convolutional backbone for frame-wise SED predictions further improves results, though its impact is smaller than that of KD.
@@ V Conclusion
In this paper, we proposed and rigorously evaluated multiple efficient architectures and model sizes for sound event detection. Our findings demonstrate that incorporating well-suited sequence models on top of a convolutional backbone consistently enhances performance compared to fully convolutional models. Additionally, the optimal sequence model choice depends on the specific complexity metric being prioritized. We further emphasized the critical role of knowledge distillation in achieving high performance and showed that, with an optimized distillation strategy, our models can match state-of-the-art transformer performance while using only around 5% of their parameters. To foster further advancements, we release our pre-trained models along with our training and evaluation framework, inviting researchers to develop new sequence models that push the performance–complexity trade-off even further.

### v2

@@ Abstract
Prior research on low-complexity audio classification has leveraged efficient convolutional neural network (CNN) architectures adapted from the vision domain [6, 7, 8] or introduced custom-designed efficient CNNs [9, 10, 11, 12, 13]. However, most of these models are optimized for clip-wise predictions, i.e., predicting a class label for an entire given audio clip. In contrast, the goal in SED is to identify specific acoustic events in a recording, including begin and end times, which implies a need for frame-wise predictions [14].
Unlike clip-wise tasks, where local features from the CNN are typically aggregated using global pooling, SED commonly employs a sequence model on top of the CNN to capture long-term temporal dependencies. Efficient SED architectures frequently utilize depthwise separable convolutions to reduce complexity [15, 16, 17], yet the optimal choice of sequence model remains an open question. While recurrent neural networks (RNNs) are a widely adopted default—forming the basis of the popular convolutional recurrent neural network (CRNN) architecture [18, 19, 20, 21]—recent research has also shown promise in lightweight attention-based models [15, 17].
The present paper addresses this gap by systematically comparing various sequence models built upon a fixed CNN backbone, aiming to identify sequence models that offer the best trade-off between performance and complexity. We choose MobileNetV3 [22] as the CNN backbone, as it is one of the most widely used depthwise separable CNNs and has demonstrated strong performance in large-scale audio tagging [8]. By fixing the CNN architecture, we focus our analysis entirely on the temporal modeling stage.
To get the most meaningful results, we evaluate on the largest publicly-available SED dataset - AudioSet Strong [23] - and use the state-of-the-art training pipeline [4], which includes knowledge distillation [24] from large transformer models. Our experiments aim to answer the following key questions:
@@ II Architectures
where and represent a linear projection aligning embeddings with the dimension of the sequence model . If no sequence model is used, and the linear projection (, ) are identity functions. The output of is then transformed by a position-wise linear layer, parameterized by and , to generate frame-level predictions , with dimension corresponding to the number of classes in the task.
@@ II-A Frame-wise Convolutional Backbone
As the convolutional backbone, we use MobileNetV3 [22]. To enable frame-wise instead of clip-wise predictions, we remove the global pooling and the global classification heads. The outputs of this network are then three-dimensional embeddings of shape . Compared to the original MobileNet, we increase some convolutional strides along the frequency axis from 1 to 2 while decreasing others along the time axis from 2 to 1. This changes the shape of from to , matching the temporal resolution of the pseudo-labels used for knowledge distillation as described in Section III-B. Therefore, each of the 250 time frames in corresponds to a 40-millisecond segment of the original 10-second audio clip. We refer to this adapted model as frame-wise MobileNet (fmn).
For the experiments that require it, we scale the frame-wise MobileNets following the method described in [8], maintaining the number of layers while adjusting the model’s width (by multiplying the default number of channels of all convolutional layers) using a scaling factor . We consider models with which we name fmn04, fmn06, fmn10, fmn20 and fmn30, respectively.
@@ III-B Training Configuration
We initialize the backbone CNNs with ImageNet [35] pre-trained weights and train them on AudioSet Weak [36], following the knowledge distillation routine from [8]. However, we find that extending the training to 300 epochs improves performance.
Following the AudioSet Weak pre-training, the full models (pre-trained backbone + sequence model) are trained on AudioSet Strong for 120 epochs using an Adam optimizer with no weight decay, a cosine learning rate schedule with 5,000 warmup steps, and a batch size of 256.
@@ IV-A Performance-Complexity Trade-Off of Sequence Models
Figure 1 shows that all sequence models improve the performance of fmn10, even with the smallest hidden dimension. However, only models incorporating TF, BiGRU, ATT, and HYBRID achieve performance comparable to or surpassing fmn20 while maintaining lower parameter counts, lower MACs, and increased throughput, resulting in a more favorable performance-complexity trade-off. This shows that adding suitable sequence models on top of fmn10 scales more efficiently than simply scaling up the convolutional backbone itself. In contrast, MAMBA and TCN fail to reach the performance level of fmn20 at any hidden dimension size.
The choice of the model offering the best performance-complexity trade-off depends on the complexity metric of interest. The top choices for parameters, throughput, and MACs are HYBRID, TF, and BiGRU, respectively. For example, fmn10 using TF and a hidden dimension of 256 (fmn10+TF:256) matches the performance of fmn20 while achieving 1.76 times its throughput and requiring only 37.6 of its total parameters. Similarly, fmn10 with HYBRID at a hidden dimension of 128 (fmn10+HYBRID:128) has fewer parameters than fmn10 without a sequence model444 If the down-projection, sequence model, and linear head together require fewer parameters or MACs than a single large linear layer mapping embeddings to target classes, adding a small sequence model can be more efficient than using no sequence model with respect to these metrics. yet improves the PSDS1 score from 42.28 to 43.45. However, fmn10+HYBRID:128 exhibits lower throughput compared to fmn10+TF:256. In terms of MACs, BiGRU is the most effective555BiGRUs involve multiple operations, such as gating, point-wise products, and point-wise summations, which do not count as MACs., but its sequential processing limits parallel computation, resulting in lower throughput compared to TF and HYBRID. We conclude that TF is the most balanced choice with strong results on all three complexity measures. In particular, fmn10+TF:256 strikes an excellent trade-off between performance gain and minimal added complexity, making it the preferred model for further experiments.
@@ IV-B Scaling the Backbone and Sequence Model
Three key conclusions can be drawn from Figure 2: 1) TF and HYBRID consistently outperform fmns without sequence models across all scaling factors; 2) HYBRID performs best in the lower complexity range, with fmn04+HYBRID and fmn06+HYBRID surpassing their TF and BiGRU counterparts; and 3) TF and HYBRID exhibit greater stability than BiGRU, as the performance drop of BiGRU from to is a consistent trend across multiple independent runs.
@@ IV-C Impact of KD and CNN Backbone Fine-Tuning
| Model | KD | Frozen CNN Backbone | PSDS1 |
@@
As outlined in Section III-B, we first pre-train the CNN backbone on AudioSet Weak and fine-tune the full model (backbone + sequence model) on AudioSet Strong using KD. Table I presents the resulting performance effects of this setup. Regardless of whether a sequence model (TF:256 on top of fmn10) is used, KD significantly enhances performance. Similarly, fine-tuning the convolutional backbone for frame-wise SED predictions further improves results, though its impact is smaller than that of KD.
@@ V Conclusion
In this paper, we proposed and thoroughly evaluated efficient sound event detection (SED) architectures. We utilized modified MobileNetV3 models, adjusted to various complexity levels, as the convolutional backbone, and combined them with different sequence models. Our findings demonstrate that pairing appropriate sequence models with the backbone consistently outperforms fully convolutional versions across all complexity levels tested. The optimal sequence model depends on the prioritized complexity metric. We further emphasized the critical role of knowledge distillation in achieving high performance and showed that, with an optimized distillation strategy, our models can match state-of-the-art transformer performance while using only around 5% of their parameters. To aid future work, we release our pre-trained models along with our training and evaluation framework, inviting researchers to develop new sequence models that push the performance–complexity trade-off even further.